The length of a rectangle is 3m less than twice the width. the perimeter of the rectangle is 42m. find the length and the width of the rectangle
Answer by Paul(988)

Length = 2x-3
Width = x
Perimenter = 42
Formula = 2(L+W)=P
2(2x-3+x)=42
2(3x-3)=42
6x-6=42
6x=48
x=8
2(8)-3=13

Hence the width is 6m and the length is 9m.
Paul.